a sample of carbon dioxide gas has a volume of 5.00 L at 35 degree celcius. What is the volume in L, will the gas occupy at 125 degree celcius?
Answer #1

Given:
Vi = 5.00 L
Ti = 35.0 oC
= (35.0+273) K
= 308 K
Tf = 125.0 oC
= (125.0+273) K
= 398 K

use:
Vi/Ti = Vf/Tf
5.00 L / 308.0 K = Vf / 398.0 K
Vf = 6.461 L
Answer: 6.46 L